an acid solution was made by mixing 1L of a 40% acid solution and 4L of a 70% solution. what is the concentration of the mixture? please put steps on how you solved it
Answer by stanbon(75887)

an acid solution was made by mixing 1L of a 40% acid solution and 4L of a 70% solution. what is the concentration of the mixture?
------
Equation:
acid + acid = acid
----
1*0.40 +4*0.70 = x*(1+4)
Multiply thru by 100 to get rid of the decimals:
------
40 + 70*4 = 500x
320 = 500x
x = 0.64 = 64%
==================
Cheers,
Stan H.